OpenStudy (anonymous):

Two lines are perpendicular. If one line has a slope of -1/13, what is the slope of the other line? answers A. -13 B. 1/13 C. 13 D. -1/13

ganeshie8 (ganeshie8):

Hint : product of slopes of perpendicular lines is -1

OpenStudy (anonymous):

So it would be the same?

ganeshie8 (ganeshie8):

nope, suppose the required slope is \(m\). since the product of slopes is -1 we have : \[m \left(\frac{-1}{13}\right) = -1\] solve \(m\)

OpenStudy (anonymous):

it would be a positive 13

OpenStudy (anonymous):

So my answer would be C.

OpenStudy (anonymous):

correct?

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

so would i do that formula for every problem like that?

ganeshie8 (ganeshie8):

Yes you may memorize that the slopes of perpendicular lines multiply to -1 :)

OpenStudy (anonymous):

ok what if its a positive? would it still be m(1/5)=-1? @ganeshie8

ganeshie8 (ganeshie8):

are you given slope of one line and asked to find the slope of a perpendicular line ?

ganeshie8 (ganeshie8):

if so, you're right ! just solve m..

OpenStudy (anonymous):

The lines below are parellel if the slope of the green line is 1/5 what is the slope of the red line? Its the same question but with a positive slope.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

m(1/5)=-1 = -5 right?

ganeshie8 (ganeshie8):

could you attach the diagram ?

ganeshie8 (ganeshie8):

only the slopes of perpendicular lines multiply to -1. not every pair of lines.. i need to see the red and green lines before confirming. .

OpenStudy (anonymous):